# Measurement-induced physics and stochastic dynamics in systems of single-body quasiparticle and many-body microscopic cases

### v2

@@ IV.3 Realizations under measurement
V Stochastic evolution with Markovian and non-Markovian dynamics
@@
For non-Markovian dynamics non-local in time, the dynamics can nolonger be described by Lindblad master equation and the corresponding stochastic Schrodinger equations can describes only the no-jump process (controlled dissipation), instead of the jump (unconditioned dissipations) upon a measurement record[34]. There is finite detector bandwidth and noise-noise correlation due to the information backflows from environment to system (in contrast to the Markovian system with nonunitary Lindblad dynamics[116]), despite the description by Markovian master equation still hold on average[34] A vanishingly small dissipative incoherent case $\gamma\rightarrow 0$ is applied to non-Markovian dynamics to guarantees the fast relaxation and dependence on noise correlation time, in the absence of quantum jump [90, 80, 96, 102]. Thus, despite the coupling to the non-Markovian environments, the continuous measurement by homodyne or heterodyne detection can be realized in $sq$-case in Markovian limit (i.e., infinite bandwidth, infinite count rate, and perfect detector), where the large dissipation rate guarantees the thermal state and the loss of initial information due to the environment. In $sq$-case, the volume-law (saturated) entanglement in short-time (long-time) scale is dominated by the Markovian (non-Markovian) dynamics.
Differently, in $mm$-case, evolution of the system strongly coupled to a thermal reservoirs (Markovian environment) is well described by the Markovian Lindblad master equation, and the non-Markovian memory effect induced complex coherence pattern (the infinite bound states due to the dissipationless non-Markovian dynamics), whose backaction effect induce the oscillation in entanglement, and the steady state is possible only in long-time limit. Note that despite the strong system-bath coupling, the Markovian thermal environment equilibrium at a certain temperature initially coherent with the system (without the quantum correlation/entanglement) in a product form $\rho_{S}\otimes\rho_{E}$[102, 103]. In this case, the strong noise in Markovian (noisy) environments results in a long timescale of slow relaxation, and the long non-Markovian noise-noise correlation time (weak non-Markovian) within bath in fact has a indistinguishable effect with the Markovian white noise, where both give rise to the classical diffusion in the system.
In interacting case, the dissipative interaction is represented by a decoherence matrix which push the system toward thermalization, In noninteracting limit, ${\bm{\gamma}}_{0}=\gamma{\bf I}$ while in the all-to-all interacting limit (Dicke limit in canonical setting[33, 32]), ${\bm{\gamma}}_{D}=|\sqrt{\gamma}\rangle\langle\sqrt{\gamma}|$, with $|\sqrt{\gamma}\rangle$ the vector consist only the element $\sqrt{\gamma}$. The vectors made of eigenvalues arranged in descending order for ${\bm{\gamma}}_{D}$ supermajorize that of ${\bm{\gamma}}_{0}$.
@@
where the expectation (or ensemble average) on the Poisson random stochastic increment $dN_{i}(t)$ is the Born probability for unit detection result, and is unknown until the measurements are done, which reads ${\mathbb{E}[dN_{i}(t)]}=\langle\psi(t)|L^{{\dagger}}_{i}L_{i}|\psi(t)\rangle\gamma_{i}dt=||L_{i}|\psi(t)\rangle||^{2}\gamma_{i}dt$, $dN_{i}(t)dN_{j}(t)=\delta_{ij}dN_{i}(t)$. This makes it unable to eliminate or multiplicate the noise[91, 45], and meanwhile, retains the Poissonian randomness, in constract to the Wiener process with Ito stochasticity.
The local (short-range) correlations in single-particle picture (closed system) corresponds to the quantity linear on the stochastic quantum states like the expectation in basis of stochastic states, and the absence of long-range order is consistent with the result of cluster decomposition[79, 80].

### v4

@@ IV.3 Realizations under measurement
Figure 1: Von Neumann entropy $-{\rm ln}{\rm Tr}D^{2}$ for $sq$ (left) and $mm$ (right) cases, which can be termed as microscopically defined entropy and thermodynamic entropy, respectively. Blue, orange to green dots represent the increasing system size at balanced partitions. The transition of linear (volume law) scaling to saturation toward a system-size-independent value (area law) can be seem in $sq$-case. The linear growth of entropy in early time is consistent with the scrambling (spreading of local information) with constant slope described by classical Lyapunov exponent[160], Thus for different system size, there are different scrambling time but same Lyapunov exponent. while the quantum mechanics with interacting effect only sets in latter as exhibited by the power-law growth toward the maximal entropy which corresponds to the dip in purity (see Fig.7). Note that the maximal entropy here does not corresponds to the maximal mixture as shown in Fig.7. The entropy in $mm$-case is dominated by the quantum mechanics. The revival behavior in $mm$-case results in a periodically fluctuating entropy, which should be distinguished from the macroscopic mechanical oscillations[180] and can be fitted by the conformal field theory (discontinuous) in the analysis of finite size scaling[39, 149, 166]. We also verify that the entropy here equals to the negative of the logarithmic negativity $-{\rm ln}({\rm Tr}D^{1/2})^{2}$[155]. The scheme (a) automatically arrives the balanced partition due to the densest measurement; While the scheme (b) admits arbitary partitions as revealed by the entanglement between the unitarily or nonunitarily evolved pure state $|\rho_{0}\rangle\langle\rho_{0}|$ and the rest. For (a), it is reasonable to suspect that, the peak of entropy between the (sub)volume phase in early time and area law phase in late time, which sharpens with increasing size, will vanish in the large system size limit, and the scrambling will also vanish.
Figure 2: Evolutions of trace and norm of the dissipative term for $sq$ (blue) and $mm$ (orange) cases.. The exponential decay and incoherence feature can be seem from the $sq$-case, reflecting that the dissipative terms is nondegenerated or nondefectively degenerated. While the algebraic decay is hard to recognized from the $mm$-case, due to the complicated coherence (averaged) pattern. The left panel is indeed a conditioned expectation of the stochastic object which does not depend on Wiener increment (Markovian noise) but on stochastic increments at earlier times[119].
Figure 4: (a) and (b) show the $e^{i(\rho_{i}+\rho_{j})t}-e^{i\rho_{i}t}e^{i\rho_{j}t}$ ($\neq 0$ only for $\rho^{{\dagger}}=\rho$) representing the periodically oscilatting classical correlations (or noncommutativity) between the two arbitary states of the $mm$-case, for incomplete and complete basis, respectively. Such pairwise correlations is unique to the $mm$-case, due to the lacking of macroscopic oscillation[180]. Blue, orange, and green dots represent the real, imaginary, and absolute parts. Such correlations, which exist between arbitarily two components in the $mm$-case, cause the delocalization and the weakened boundary driving effect. This also make the construction of Floquet driving potential available, and such correlation (in constrast to the entanglement) can also be reflected by the separability across all bipartitions[180, 182], which is realized by the projective measurements on one subsystem and cause the statistical mixture of pure states (chaotic projected ensemble) in the another subsystem (disentangling)[181]. Coherences are allowed between these mixed subsystems. (c) and (d) show the trace distace $\frac{1}{2}||D_{i}-D_{j}||$ (|||||$\cdot$||| | ⋅ | | representing the trace norm), which witness the distinguishability between states $D_{i}$ and $D_{j}$ (initially pure states), for incomplete and complete basis, respectively, with blue and orange dots correspond to the $sq$-case (blue) and $mm$-case (orange), respectively. For complete basis, a uncontinuously reducution of trace distance for $sq$-case reveals that the existence of non-Markovian process in early stage. Despite not being shown here, we have verified that, the Jensen-Shannon divergence[179] of $sq$-case reduces to zero at long-time. Still, the non-Markovianity in $mm$-case, as revealed by the memory effect and environment-to-system information backflow, causes the nonmonotonic behavior of trace distance as well as the distinguishability between the two quantum states, and a seperable mode into the hot and cool ones. Similar nonmonotonic behavior in entanglement and decoherence function are studied in Refs.[178, 183]. The trace distance in (d) for $mm$-case exhibits persistent temporal fluctuations despite more and more homogeneous compares to initial as indicated by the dashed line. It is the correlations cause the revival of trace distance and environment-to-system backflow of information. Ref.[173] indicates a possible realization of such a scenario by the balanced interplay between interaction (with bath) and tunneling.
Figure 5: Von Neumann entropy for product states $(\sum_{i}|\psi_{i}\rangle)(\sum_{i}\langle\psi_{i}^{*}|)$ ($sq$) and $(\sum_{i}|\psi_{i}\rangle)(\sum_{i}\langle\psi_{i}|)$ ($mm$). The fixed bulk pattern here indicates the possibility to be free from thermaliation, which disappears in $sq$-case with the nonunitary evolution which varies the upper and lower boundaries. Similar to Fig.1(a), a gapless (volume law with density depinning) to gapped (area law with density pinning) transition can be seem for $sq$-case (i.e., nonequilibrium boundary to equilibrium boundary crossover[157] with exponentially slow disentangling consisting with Lorentzian distribution). For product state in $mm$-case, even the local thermalization (induced by the coupling to degrees of freedom of bath) cannot be seen. The fully engineered entropy evolution in long-time (by the dissipative environment) in $sq$-case reflects the difference between the few-body cluster entanglement and the many-body (fluctuating) one.
Figure 6: (Left) $|{\rm Tr}[\frac{D_{mm}(t)}{||D_{mm}(t)||}]|$ (blue), $|{\rm Tr}[\frac{e^{-i\rho_{mm}t}}{N||e^{-i\rho_{mm}t}||}]|$ (orange), $|{\rm Tr}[\frac{D_{sq}(t)}{||D_{sq}(t)||}]|$ (green), and $|{\rm Tr}[\frac{e^{-i\rho_{sq}t}}{N||e^{-i\rho_{sq}t}||}]|$ (carmine). The last two approach to $\sqrt{\frac{1}{N}}$ and $\frac{1}{N}$, respectively, in long-time. The variance of non-Hermitia observable[180] as well as the Jensen-Shannon divergence (JSD) also support the conclusion that the $mm$-case has much robust uncertainty compares to the $sq$-case. Note that here the uncertainty and randomness are not due to the selections by measurement apparatus, which mean the Green function here does not contains a self-energy term. (Right) The expectation $|{\rm Tr}[\frac{D_{sq}(t)}{||D_{sq}(t)||}]|=\langle\rho_{0}|e^{i\rho_{sq}t}|\rho_{0}\rangle$ (Orange; without consider the perturbation from measurement apparatus) and $\langle\rho^{\prime}_{0}|e^{i\rho_{sq}t}|\rho^{\prime}_{0}\rangle$ (Blue; consider the perturbation from measurement apparatus). The effect measurement apparatus here is small but finite due to the nonunitarility of $e^{i\rho_{sq}t}$. The traditional mass perturbation theory is invalid here. The postselection probability after measurement $|\langle\rho^{\prime}_{0}|\rho_{0}\rangle|^{2}$ (in the absence of interaction) is shown in color green. While in $mm$-case, the projective measurement suppress the long-range entanglement but does not cause the selections on measured states due to the effect of unique Markovian bath. Another reason is that we choose a time-independent basis $\langle\rho_{0}|\rho_{0}\rangle$. Consequently, the mapping $|\rho_{0}\rangle\rightarrow|\rho^{\prime}_{0}\rangle$ does not cause any new conservations. This indeed reflects the existence of response to the global quench in $mm$-case, just like the complete thermal state in late stage of $sq$-case, i.e., the common characteristic between the Markovian system in $sq$-case and the unique Markovian bath in $mm$-case.
Figure 7: Purities ${\rm Tr}|D|^{2}$ for $D=D_{sq}$ (a) and $D=D_{mm}$ (b). Despite the global symmetry in $sq$-case, the purification back to the pure state is only available in the $mm$-case. This is because, in $sq$-case, the non-Markovian bath plays an essential role in engineering the dynamics of the system. The inset in (a) shows the entropy (the same with Fig.1(a)) for a comparasion. There are two times of entanglement enhancement (purity decaying) in $sq$-case, which are due to the scrambling and the combination of monitoring and boundary-driven effect, respectively. The red arrows in (a) indicate the time when the long-range correlation as well as the area-law scaling become dominating (entanglement becomes nonextensive due to the absence of scrambling), implying the exponential decay and strong localization[159] such that the Gaussian (exponentiated bilinear) is achievable. In late-stage of $sq$-case (after the red arrows), the continuous monitorng from the dissipative environment cause the decay of entropy to the finite value.
V Method and Numerics
We define the densities of the two classes as $\rho_{mm}:=\sum_{i}|\psi_{i}\rangle\langle\psi_{i}|$, $\rho_{sq}:=\sum_{i}|\psi_{i}\rangle\langle\psi^{*}_{i}|$. For $\rho_{sq}$ such asymmetry construction (off-diagonal) of the density is also consistent with that in the content of Dicke states[169, 171]. The reduced form can be seem by the state of the combined system environment $|\rho_{SE}\rangle=\sum_{i}|\psi_{i}\rangle|\mathcal{E}_{i}\rangle$ with $\langle\mathcal{E}_{i}|\mathcal{E}_{j}\rangle=\langle\mathcal{E}_{i}|\mathcal{E}_{j}^{*}\rangle=\delta_{ij}$. The complex eigenstates are prepared aimed to obtaining the non-Hermitian operator ${\bf\Psi}$ whose elements follow the distribution of Gaussian ensemble, upto second and fourth moments, and reads, when consider two real eigenvalues, ${\bf\Psi}=\lambda_{\alpha}{\bf\Psi}_{\alpha}+\lambda_{\beta}{\bf\Psi}_{\beta}$ For a general route of construction, see Refs.[150, 8]. The densities $D_{mm/sq}(t)=e^{i\rho_{mm/sq}t}(\sum_{i}|\psi_{i}\rangle)(\sum_{i}\langle\psi_{i}|)$, whose square of trace is the Loschmidt echo, is important in revealing the difference between $sq$ and $mm$ cases. The initial state for both cases are pure state and unentangled. Available convex sum form for densities in $mm$-case (diagonal in the Schmidt basis[156]) agrees with the absence of topological order therein. Environmental decoherence effect (off-diagonal fluctuation) emerges when $\langle\mathcal{E}_{i}|\mathcal{E}_{j}\rangle\neq 0$ for $i\neq j$, and this will suppress the entanglement and cause the gapless structure for $mm$-case and early stage of $sq$-case. Such decoherence and off-diagonal fluctuation will be suppressed by larger system size in $sq$-case, and decay with time toward the gapped (area law). Also, the entropy for product state $(\sum_{i}|\psi_{i}\rangle)(\sum_{i}\langle\psi_{i}^{*}|)$ converges to $1/2^{N}\approx 0.0039$ (maximally mixed density) where $N=8$ here is the maximal number of independent states, which indicating the possible relation to the Page value[161]. Also, as shown in Fig.4, the correlation dynamics revealed by the expectation of covariance matrix (nonlocal density; fermion Green’s function)[164, 129, 170] ${\rm Tr}[D_{mm/sq}(t)]$, reflects that in $sq$-case a trajectory in a $2^{N}$-dimensional space can be encoded into the evolution of the $N\times N$ observable $e^{-i\rho_{mm/sq}t}$, where $\overline{\rho_{sq/mm}}=\overline{e^{-i\rho_{sq/mm}t}}_{t=0}=\lim_{t\rightarrow\infty}|{\rm Tr}[\frac{e^{-i\rho_{sq}t}}{N||e^{-i\rho_{sq}t}||}|=(\lim_{t\rightarrow\infty}|{\rm Tr}[\frac{D_{sq}(t)}{||D_{sq}(t)||}]|)^{2}=\frac{1}{N}$. A polynomial complexity in system size is only available in the late-stage of $sq$-case, where the simplification is available due to the absence of persistive fluctuates in Hilbert space[166] and the Lieb-Robinson bounds and nonunitary scrambling light cones. Also, the incoherent coupling to the environment is absent in the late-stage of $sq$-case which is in agree with Refs.[164, 129] . While such decoherence as well as stochastic dynamics is persistent in $mm$-case, due to the restricted Hilbert space and disorders which guarantee the accessibility of (maximally mixed) infinite-temperature featureless equilibrium state in long-time limit. Despite the classical stochastic processes, like the random diffusion, still observable in $mm$-case, the dynamics periodically back to the initial state, consistent with the microscopic theory. Such recurrence is recently studied in terms of the uncertainty relation[163, 165] and hybrid quantum circuit[166].
[…]
Consistent with the results in Ref.[184], here the non-Hermitian and unitary observables $O_{nH}=e^{-i\rho_{mm}t}$ ($t\neq 0$) is full-rank such that it can be decomposed into $O_{nH}=U_{mm}R_{mm}$ with positive definite $R_{mm}=\sqrt{O_{nH}^{{\dagger}}O_{nH}}$ and $U_{mm}=\frac{O_{nH}}{\sqrt{O^{{\dagger}}_{nH}O_{nH}}}$ ($U_{mm}^{{\dagger}}U_{mm}=I$). Then we can verify that $\langle\rho_{0}|O_{nH}|\rho_{0}\rangle=\langle\rho^{\prime}_{0}|R_{mm}|\rho_{0}\rangle=\langle\rho^{\prime}_{0}|O_{nH}|\rho^{\prime}_{0}\rangle$, with $|\rho^{\prime}_{0}\rangle=U_{mm}^{{\dagger}}|\rho_{0}\rangle$ and $\langle\rho^{\prime}_{0}|R_{mm}|\rho_{0}\rangle/\langle\rho^{\prime}_{0}|\rho_{0}\rangle=1$. The role of measurement apparatus can then be taken into account by preselecting the $|\rho_{0}\rangle$, or equivalently, postselecting the $|\rho^{\prime}_{0}\rangle$, where the perturbations from measurement can be revealed by the nonzero imaginary part ${\rm Im}\langle\rho^{\prime}_{0}|R_{mm}|\rho_{0}\rangle/\langle\rho^{\prime}_{0}|\rho_{0}\rangle\neq 0$. Thus the mapping $|\rho_{0}\rangle\rightarrow|\rho^{\prime}_{0}\rangle$ does not generates any new symmetries, instead, zero imaginary part of $\langle\rho^{\prime}_{0}|R_{mm}|\rho_{0}\rangle/\langle\rho^{\prime}_{0}|\rho_{0}\rangle$. guarantees the existence of response to the global quench in $mm$-case.
In the numerics as shown in Fig.2, we consider the case totally dominated by the dissipative process such that the unitary part (i.e., evolution and decoherence due to the system Hamiltonian) in Lindblad master equation is vanished. This cause the complete fail of perturbative treatment (e.g., the one introduced in Ref.[133] which is implemented by considering the steady state of nonperturbed part be a pure state).
An efficient way to enhance the noise from the environment’s measurement is replacing the $-{\rm ln}{\rm Tr}D^{2}$ by $-{\rm ln}\sum_{i}{\rm Tr}\rho_{i}^{2}$ where $\sum_{i}\rho_{i}=D$, which cause a long-time steady state closer to a product state with much lower entanglement compares to Fig.1 and similar in mechanism to Fig.LABEL:FK (by fill in the off-diagonal components such that it becomes fail to estimate the expectation value from local observable like the single-particle macroscopic sample). Obviously, this also corresponds to a purifying for $sq$-case.
For Loschmidt echo shown in Fig.3, we define the initial state $|\rho_{0}\rangle=\sum_{i}|\psi_{i}\rangle$, and estimate the overlap $\mathcal{L}(\rho,\rho^{\prime},\cdots):=\frac{|\langle\rho_{0}|e^{-i\rho t}e^{i\rho^{\prime}t}\cdots|\rho_{0}\rangle|^{2}}{||e^{i\rho t}e^{-i\rho^{\prime}t}\cdots||^{2}}$. Note that for $sq$-case, due to the nonunitarility, a artifical renormalizaton by the squared norm is necessary to avoid the diverging as we shown in Ref.[147]. In nondegenerated region, eigenstates in $sq$-case are mutually independent, while those in $mm$-case are mutually dependent. This can be easily verified by comparing $\mathcal{L}(\rho+\rho^{\prime},\cdots)$ and $\mathcal{L}(\rho,\rho^{\prime},\cdots)$, e.g., $\frac{|\langle\rho_{0}|e^{-i\rho_{mm}t}|\rho_{0}\rangle|^{2}}{||e^{-i\rho_{mm}t}||^{2}}$ and $\frac{|\langle\rho_{0}|e^{-i\sum_{i=1}^{j}|\psi_{i}\rangle\langle\psi_{i}|t}e^{-i\sum_{i=j+1}^{N}|\psi_{i}\rangle\langle\psi_{i}|t}|\rho_{0}\rangle|^{2}}{||e^{-i\sum_{i=1}^{j}|\psi_{i}\rangle\langle\psi_{i}|t}e^{-i\sum_{i=j+1}^{N}|\psi_{i}\rangle\langle\psi_{i}|t}||^{2}}$. While in NDD region, each eigenstate therein is independent of arbitary other one, producing the long range by system-size-independent correlations. The most significant feature of this region is the two parallel by seperated boundaries in echo spectrum. This is a special region where the gap is not supported by coherences and will not prohibit the global thermalization (when there are system-size-dependent correlations).
VI Conclusion
The steady state of $mm$-case (consistent with the open boundary condition) is only due to the skin effect, where the many-body state (without exponential localization) holds global fluctuation due to the Hermiticity and the long-time behavior can be well predicted by diagonalization. This is in contrast to the localized state in $sq$-case (consistent with the periodic boundary condition) where the single-particle eigenstates exponentially decay with the localization length independent of system size[56, 94, 10, 124, 125, 130], and thus the long-time behavior is harder to predict by approximately diagonalization due to the non-Hermiticity. Zeno effect and confined light cone[131, 96] are unique to the $sq$ case, while Hilbert space fragmentation[154, 121] and linearized bosonization are unique to the $mm$ case, and the topological islands[125] is unique to the intermedia regime with finite strength of disorder but not yet reaches the ensemble-averaged result. A similar study with the correspondence to quantum systems is available by by, e.g., transverse-field Ising model[162], XXX spin model, or BoseHubbard model[161], where the ergodicity (late stage) and scrambling (early stage) of $sq$-case are shared by the qubit ladder and nonintegrable transverse-field Ising model, while the nonergodicity of $mm$-case is shared by the qubit chain and integrable XXX spin model. Also, a protocol similar to the continuous measurement which give rise to the diffusion is the collective jumps[168] where the collectiveness can be diagnosed by the enhanced entanglement. Intermediate cases is achievable through complex disorder patterns by turning the transverse and longitudinal magnetic fields where the scrambling occurs with temporal fluctuations suppressed by the disorders (ensemble average) and initial state becomes mixed up in a large subspace[162]. While the absence of scrambling in $mm$-case is due to the conservation of initial states and each subspace (block-diagonal unitary operator) owns a unique exponentially decaying behavior, anology to the internal dynamics of black hole and local integrals of motion of many-body localization which suppress the scrambling and rise the fluctuations. Note that the constriction on Hilbert space of the $mm$-case should be distinguished from the restriction of Hilbert space in $sq$-case where for the latter it is implemented by the construction of, e.g., Dicke basis (classical mixture cause thermalization as guaranteed by the permutational symmetry) and reducing from the $O(2^{N})$ Hilbert space dimension to $O(N)$ subspaces.
Data Availability Statement: The data that support the findings of this study are available within the article.
VII Appendix: Stochastic evolution with Markovian and non-Markovian dynamics
@@
For non-Markovian dynamics non-local in time, the dynamics can nolonger be described by Lindblad master equation and the corresponding stochastic Schrodinger equations can describes only the no-jump process (controlled dissipation), instead of the jump (unconditioned dissipations) upon a measurement record[34, 167]. There is finite detector bandwidth and noise-noise correlation due to the information backflows from environment to system (in contrast to the Markovian system with nonunitary Lindblad dynamics[116]), despite the description by Markovian master equation still hold on average[34] A vanishingly small dissipative incoherent case $\gamma\rightarrow 0$ is applied to non-Markovian dynamics to guarantees the fast relaxation and dependence on noise correlation time, in the absence of quantum jump [90, 80, 96, 102]. Thus, despite the coupling to the non-Markovian environments, the continuous measurement by homodyne or heterodyne detection can be realized in $sq$-case in Markovian limit (i.e., infinite bandwidth, infinite count rate, and perfect detector), where the large dissipation rate guarantees the thermal state and the loss of initial information due to the environment. In $sq$-case, the volume-law (saturated) entanglement in short-time (long-time) scale is dominated by the Markovian (non-Markovian) dynamics.
Differently, in $mm$-case, evolution of the system strongly coupled to a thermal reservoirs (Markovian environment) is well described by the Markovian Lindblad master equation, and the non-Markovian memory effect induced complex coherence pattern (the infinite bound states due to the dissipationless non-Markovian dynamics), whose backaction effect induce the oscillation in entanglement, and the steady state is possible only in long-time limit. Note that despite the strong system-bath coupling, the Markovian thermal environment equilibrium at a certain temperature initially coherent with the system (without the quantum correlation/entanglement) in a product form $\rho_{S}\otimes\rho_{E}$[102, 103]. In this case, the strong noise in Markovian (noisy) environments results in a long timescale of slow relaxation, and the long non-Markovian noise-noise correlation time (weak non-Markovian) within bath in fact has a indistinguishable effect with the Markovian white noise, where both give rise to the classical diffusion in the system. In interacting case, the dissipative interaction is represented by a decoherence matrix which push the system toward thermalization, In noninteracting limit, ${\bm{\gamma}}_{0}=\gamma{\bf I}$ while in the all-to-all interacting limit (Dicke limit in canonical setting[33, 32, 168]), ${\bm{\gamma}}_{D}=|\sqrt{\gamma}\rangle\langle\sqrt{\gamma}|$, with $|\sqrt{\gamma}\rangle$ the vector consist only the element $\sqrt{\gamma}$. The vectors made of eigenvalues arranged in descending order for ${\bm{\gamma}}_{D}$ supermajorize that of ${\bm{\gamma}}_{0}$.
@@
where the expectation (or ensemble average) on the Poisson random stochastic increment $dN_{i}(t)$ is the Born probability for unit detection result, and is unknown until the measurements are done, which reads ${\mathbb{E}[dN_{i}(t)]}=\langle\psi(t)|L^{{\dagger}}_{i}L_{i}|\psi(t)\rangle\gamma_{i}dt=||L_{i}|\psi(t)\rangle||^{2}\gamma_{i}dt$, $dN_{i}(t)dN_{j}(t)=\delta_{ij}dN_{i}(t)$. This makes it unable to eliminate or multiplicate the noise[91, 45], and meanwhile, retains the Poissonian randomness, in constract to the Wiener process with Ito stochasticity. Up to large number limit where the homodyne scheme and quantum state diffusion are achieved, the Poisson $dN_{i}$ is replaced by the multiplicative Gaussian Ito noise. The local (short-range) correlations in single-particle picture (closed system) corresponds to the quantity linear on the stochastic quantum states like the expectation in basis of stochastic states, and the absence of long-range order is consistent with the result of cluster decomposition[79, 80].